Relation btn thermal coefficient and volumetric coefficien 
 

Grade:11

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

Please complete the question

Regards
Arun
3 years ago
α = coefficient of linear expansion
β = coefficient of areal expansion
γ = coefficient of volume expansion
Since, coefficient of linear expansion is in one direction and areal in two dimension followed by three dimension in case of volume expansion. So, to compensate that factor the relation between them will be like.
**α = β /2= γ /3**